Account Balance ($)
0 to 50,000 15%
50,000 to 75,000 $7,500 + 25% of the amount over 50,000
75,000 to 100,000 $13,750 + 34% of the amount over 75,000
100,000 to 335,000 $22,250 + 39% of the amount over 100,000
335,000 to 10,000,000 $113,900 + 34% of the amount over 335,000
10,000,000 to 15,000,000 $3,400,000 + 35% of the amount over 10,000,000
15,000,000 to 18,333,333 $5,150,000 + 38% of the amount over 15,000,000
18,333,333 and up 35%
Fiscal Year Gross Profit ($)
June, 2009 - May, 2010 261,100
June, 2011 - May, 2012 277,500
June, 2012 - May, 2013 292,000
June, 2013 - May, 2014 300,900
  1. Given the information provided here perform your own calculations to show the amount of corporate tax paid for each of the four (4) fiscal years covered in the table above.
  2. Is there a pattern? Explain the pattern and what it may suggest about the future performance of the organization. How can this conclusion advise investors and organizational decision makers?
  3. If there was a 20% reduction in the Gross Profit for fiscal year 6/12 - 5/13, how might this have changed the gross profit expectations for the 6/13 - 5/14 fiscal year? Provide rationale.
  4. Calculate the total expected taxes that would have been paid for these two fiscal years, with the 20% reduction in 6/12 - 5/13 and the resultant expected change in the 6/13 - 5/14 gross profit.
Answer #1

The corporate tax will be calculated as :

2009-2010= 22250+ 161100*39%= 85079

2011-2012= 22250+ 177500*39%= 91475

2012-2013= 22250+ 192000*39%= 97130

2013-2014= 22250+ 200900*39%= 100601

In the above calculation, it is clear that the amount of corporate tax is rising but the rate at which is rising is declining. For example: difference in the tax rate between fiscal year 2009-2010 and 2011-12 is 6396 and for the next fiscal year it is 5655 and so on. So rate of tax is of declining pattern. Gross profit is rising.

This would depict the investors to invest in the company as gross profits are rising rate of tax declining. It will be benefecial for them.

If there is 20% reduction in fiscal year 12- 13, gross profit will be 233600. From here we can see that gross profit declined from this year as compared to 2011-12. So the futire expectation for the year 2013-14 will also be declined.

expected tax for fiscal year 2012-13:

233600= 22250+ 133600*39%= 74354